OpenStudy (anonymous):

If the equation of a circle is (x + 5)^2 + (y - 7)^2 = 36, its radius is *6 *16 *36

hartnn (hartnn):

\((x-a)^2+(x-b)^2=r^2\) will have the radius as 'r' so, r^2 = 36 r=.. ?

OpenStudy (anonymous):

So it's36?

hartnn (hartnn):

r^2 = 36 and we need 'r' !

OpenStudy (anonymous):

OH. So it's 6!

hartnn (hartnn):

yes!